# Self-Supervised Frameworks for Speaker Verification via Bootstrapped Positive Sampling

### v2

@@ II-C SwAV
SwAV [caron2020SwAV] simultaneously clusters embeddings while enforcing consistency between assignments produced for the different views of the same utterance. Prototypes are learned jointly with the model, and soft assignments are produced online using the Sinkhorn-Knopp algorithm [cuturi2013Sinkhorn]. The “swapped” prediction mechanism implies that a view’s code is predicted from another view’s representation. Additionally, a queue $\bm{Q}^{\text{SwAV}}$ of latest embeddings is used to simulate a larger batch size and to provide more stable cluster assignments.
@@
$\mathcal{L}_{\text{SwAV}}=-\frac{1}{B\cdot P}\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}\sum_{k=1}^{P}\bm{a}_{i,k}\log\frac{\ell\left(\bm{z}_{i}^{\prime},\bm{p}_{k}\right)}{\sum\limits_{k^{\prime}=1}^{P}\ell\left(\bm{z}_{i}^{\prime},\bm{p}_{k^{\prime}}\right)},$ (3)
@@ II-D VICReg
VICReg [bardes2022VICReg] is based on three training components that maximize information directly from the embeddings and provide an explicit solution to avoid collapse. The variance ($v$) component enforces the variance to reach $1$ along the $D_{\text{emb}}$ dimensions of the embeddings to produce class-dependent representations and prevent a collapsing solution, such that $v\left(\bm{Z}\right)=\frac{1}{D_{\text{emb}}}\sum_{d=1}^{D_{\text{emb}}}\max\left(0,1-\sqrt{\operatorname{Var}(\bm{z}^{d})}\right)$. The invariance ($s$) component minimizes the distance between the anchor and positive to learn invariance to perturbations such as noise and reverberation, such that $s\left(\bm{Z},\bm{Z}^{\prime}\right)=\frac{1}{B}\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}\left\|\bm{z}_{i}-\bm{z}_{i}^{\prime}\right\|_{2}^{2}$. The covariance ($c$) component makes the off-diagonal coefficients of the embeddings covariance matrix $C$ to be close to 00 to prevent the different dimensions of the embeddings from encoding similar information, such that $c\left(\bm{Z}\right)=\frac{1}{D_{\text{emb}}}\sum_{d=1}^{D_{\text{emb}}}\sum_{\begin{subarray}{c}d^{\prime}=1\\ d^{\prime}\neq d\end{subarray}}^{D_{\text{emb}}}[C(\bm{Z})]_{d,d^{\prime}}^{2}$.
@@ II-E DINO
DINO [caron2021DINO] is a self-distillation framework based on the knowledge distillation paradigm where a student network is trained to predict the representations of a teacher network. Following the “multi-crop” strategy [caron2020SwAV], a larger set of augmented utterances on different frame lengths is considered, resulting in four small (local) and two large (global) frames. All views are fed through the student network, while only global views are fed through the teacher network to learn “local-to-global” correspondences. The teacher network is updated with an EMA of the student weights following the asymmetrical joint-embedding architecture. The student and teacher heads output a $D_{\text{emb}}$-dimensional feature normalized with a temperature-softmax. To avoid collapse, sharpening and centering are applied to the teacher outputs: centering prevents one dimension from prevailing while sharpening discourages collapse to the uniform distribution.
@@
$\mathcal{L}_{\text{DINO}}=\frac{1}{B}\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}\sum_{t=1}^{2}\sum_{\begin{subarray}{c}s=1\\ s\neq t\end{subarray}}^{6}H\bigg{(}\frac{\bm{z}_{i,t}^{\prime}-\bm{c}}{\tau_{\text{t}}},\frac{\bm{z}_{i,s}}{\tau_{\text{s}}}\bigg{)},$ (5)
where $H\left(\bm{a},\bm{b}\right)=-\operatorname{softmax}(\bm{a})\log\left(\operatorname{softmax}(\bm{b})\right)$, $\bm{z}_{i,t}^{\prime}$ is the $t$-th teacher embedding of the $i$-th sample, $\bm{z}_{i,s}$ is the $s$-th student embedding of the $i$-th sample, $\tau_{\text{t}}$ is the teacher temperature, $\tau_{\text{s}}$ is the student temperature, and $\bm{c}$ is a running mean on the teacher embeddings.
Figure 3: SSL training framework for SV with SSPS-Clustering. Compared to the standard framework, the positive is retrieved from a memory queue based on the anchor’s clustering assignment. Clustering is performed at the beginning of each epoch on reference representations computed from longer and non-augmented audio segments.
@@ III SSPS: Self-Supervised Positive Sampling
Different studies have demonstrated theoretically that the primary factors influencing the quality of SSL representations lie in how the positives are defined [balestriero2023CookbookSSL]. This incorporated “a priori” knowledge helps model class distribution, which is fundamental to learning meaningful representations without human supervision.
@@ III-A Motivation
SSL frameworks rely on data-augmentation techniques to produce an appropriate positive for an anchor. However, this approach has inherent limitations since data-augmentation highly depends on the input data and the downstream task. Moreover, standard techniques may not be sufficient to simulate the diversity among samples from the same class.
In speaker recognition, the challenge is designing data-augmentation replicating various acoustic conditions to prevent the model from relying on channel characteristics. SSL models trained for speaker verification are prone to learn channel-related information as anchor-positive pairs are sampled from the same utterances (recordings). This effect is amplified as the VoxCeleb dataset was collected ”in the wild”, implying that speech segments are corrupted with various channel effects and other factors. Conventional audio augmentation techniques, such as background noise addition and reverberation, are fundamental but fall short of limiting the impact of the same-utterance positive sampling. Therefore, SSL frameworks lack the ability to match two utterances with different channel characteristics to the same speaker identity. This represents the main bottleneck to reaching the performance of supervised systems, which inherently address this challenge using human-annotated labels.
@@ III-B Method
Self-Supervised Positive Sampling (SSPS) is proposed to sample more relevant positives using the acquired knowledge of the SSL model. Speaker recognition datasets, such as VoxCeleb, are composed of multiple recordings (or sessions) representing several utterances from a unique speaker, and each speaker has multiple recordings (at least two).
Based on this observation and the fact that SSL representations encode channel information, the following assumption is formulated: “SSL same-utterance positive sampling group elements of the same recordings, sharing similar channel information, before modeling speaker identities”.
Suppose that representations $\bm{y}_{1}$ is from speaker $S$ and recording $R$, $\bm{y}_{2}$ is from the same recording but a distinct utterance, $\bm{y}_{3}$ is from the same speaker but a different recording, and $\bm{y}_{4}$ is from a different speaker. In this case, the assumption can be expressed as $\operatorname{sim}\left(\bm{y}_{1},\bm{y}_{2}\right)\geq\operatorname{sim}\left(\bm{y}_{1},\bm{y}_{3}\right)\geq\operatorname{sim}\left(\bm{y}_{1},\bm{y}_{4}\right)$, where $\operatorname{sim}(\bm{u},\bm{v})$ is the cosine similarity between two vectors. This means that the latent space is first organized by groups of same-recording utterances and then groups of same-speaker utterances.
SSPS aims to sample accurate positives from the same speakers but from different recordings. This approach falls within the family of bootstrapped techniques, as the selection of positives relies entirely on the model’s own evolving representations, without using any labels, and improves over iterations (i.e., enhanced representations lead to better sampling and vice versa). This method can be implemented into any SSL framework and introduces a novel axis of freedom, namely, the design of the positive sampling algorithm. It extends SSL beyond tweaking the architecture of the model, the data-augmentation, or the objective function.
Positive sampling and data-augmentation are complementary strategies that serve the common goal of generating meaningful anchor-positive pairs to shape the underlying speaker distribution in the latent space. Data-augmentation is crucial when employing the standard SSL same-utterance positive sampling, since both the anchor and the positive originate from the same recording and would otherwise have very similar acoustic conditions. However, since SSPS creates anchor-positive pairs from different recordings of the same speaker, it inherently introduces greater variability, and thus has the notable property of being less reliant on data-augmentation, particularly when the training data exhibits diverse acoustic conditions. This property is demonstrated through experiments presented at the end of the article.
@@ III-C Framework
Given a training utterance $u_{i}$ with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}$, from which the anchor is sampled, let $pos(i)$ denote the index of an utterance $u_{pos(i)}$ used to sample the positive. Standard SSL creates the positive from the same utterance by setting $pos(i)=i$ (Figure 1a) and applying random data-augmentation. In contrast, SSPS selects a pseudo-positive from a different utterance in latent space, such that $pos(i)\neq i$ (Figure 1b).
The proposed framework is built around the following main components: (1) reference representations, extracted from longer unaltered audio segments to obtain a robust latent space of training samples; (2) two new memory queues, storing all reference representations and latest positive embeddings throughout training; and (3) pseudo-positive sampling, which determines $pos(i)$ based on the latent space of reference representations and retrieves the corresponding positive embedding from the memory queue.
@@ III-C1 Reference representations
Since SSPS aims to capture the underlying data patterns for identifying speakers and recordings, the anchor $\bm{y}_{i}$ and positive $\bm{y}_{i}^{\prime}$ representations are not used to determine $pos(i)$. Instead, a reference frame $\hat{\bm{x}_{i}}$, extracted from $u_{i}$ but without data-augmentation and using a longer audio segment, is introduced. As shown in Fig. 3, the corresponding reference representation $\hat{\bm{y}_{i}}$ is computed during training and used by the SSPS sampling algorithm.
@@ III-C2 Memory queues
SSPS relies on two memory queues, $\hat{\bm{Q}}$ of size $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\hat{\bm{Q}}|},D_{\text{repr}})$ and $\bm{Q}^{\prime}$ of size $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\bm{Q}^{\prime}|},D_{\text{emb}})$. $\hat{\bm{Q}}$ is used to store the reference representations $\{\hat{\bm{y}}_{i}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{I}}}$ and $\bm{Q}^{\prime}$ is used to store the positive embeddings $\{\bm{z}^{\prime}_{i}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{I}}}$. The former helps the sampling algorithm to determine which samples to use as pseudo-positives (i.e., determining $pos(i)$), while the latter is used to extract the corresponding positive embeddings for the SSL objective function. By default, $|\hat{\bm{Q}}|$ and $|\bm{Q}^{\prime}|$ are set to $N$ to cover all training samples, but $|\bm{Q}^{\prime}|$ can be reduced depending on the setup to limit memory usage. Memory queues are updated dynamically at each training iteration using $\hat{\bm{Y}}$ and $\bm{Z}^{\prime}$, respectively, as illustrated in Fig. 3. For clarity, queues are indexed by training index, meaning $\hat{\bm{q}}_{i}$ and $\bm{q}^{\prime}_{i}$ retrieve the element corresponding to the $i$-th training sample, not the $i$-th position in the queue.
@@ III-C3 Pseudo-positive sampling
After a pre-defined number of training epochs with the standard SSL framework, it is assumed that same-speaker representations are grouped in the latent space but with a large scatter due to the strong representation of channel and recording conditions. This assumption is used to sample a pseudo-positive from the neighborhood of the anchor that does not share the same channel and recording conditions. Pseudo-positives embeddings will be sampled from $\bm{Q}^{\prime}$ if and only if $pos(i)\neq\varnothing$, otherwise default SSL positive sampling is applied. For most SSL frameworks, the positive embedding of the anchor-positive pair of the $i$-th sample will be $\bm{q}^{\prime}_{pos(i)}$ extracted from $\bm{Q}^{\prime}$, instead of $\bm{z}_{i}^{\prime}$ from the mini-batch, in the objective functions $\mathcal{L}_{\text{SimCLR}}$ (1), $\mathcal{L}_{\text{MoCo}}$ (2), $\mathcal{L}_{\text{SwAV}}$ (3), and $\mathcal{L}_{\text{VICReg}}$ (II-D). For the DINO framework, student outputs are considered to be anchors and teacher outputs to be positives, which implies that the teacher embeddings of the $i$-th sample will be $\bm{q}^{\prime}_{pos(i),1}$ and $\bm{q}^{\prime}_{pos(i),2}$, instead of $\bm{z}_{i,1}^{\prime}$ and $\bm{z}_{i,2}^{\prime}$ from the mini-batch, in the objective function $\mathcal{L}_{\text{DINO}}$ (5).
Note that SSPS can be made symmetric by using $\bm{q}_{pos(i)}$, extracted from a new memory queue $\bm{Q}$ similar to $\bm{Q}^{\prime}$, as a positive for $\bm{z}_{i}^{\prime}$. However, this technique was not considered as preliminary experiments demonstrated that it results in equivalent downstream performance.

### v3

@@ II-C SwAV
SwAV [9] simultaneously clusters embeddings while enforcing consistency between assignments produced for the anchor and the positive. Prototypes are learned jointly with the model and soft assignments are generated online using the Sinkhorn-Knopp algorithm [32]. Following a “swapped” prediction mechanism, the code of the anchor is predicted from the embedding of the positive, and vice versa. A queue $\bm{Q}^{\text{SwAV}}$ of latest embeddings is used to simulate a larger batch size, thereby providing more stable cluster assignments.
@@
$\mathcal{L}_{\text{SwAV}}=-\frac{1}{B\cdot P}\sum_{i\in\mathcal{B}}\sum_{k=1}^{P}\bm{a}_{i,k}\log\frac{\ell\left(\bm{z}_{i}^{\prime},\bm{p}_{k}\right)}{\sum\limits_{k^{\prime}=1}^{P}\ell\left(\bm{z}_{i}^{\prime},\bm{p}_{k^{\prime}}\right)},$ (3)
@@ II-D VICReg
VICReg [11] maximizes information directly from the embeddings and provides an explicit solution to avoid collapse. The variance ($v$) component enforces the variance to reach $1$ along the $D_{\text{emb}}$ dimensions of the embeddings to produce class-dependent representations and prevent a collapsing solution, such that $v\left(\bm{Z}\right)=\frac{1}{D_{\text{emb}}}\sum_{d=1}^{D_{\text{emb}}}\max\left(0,1-\sqrt{\operatorname{Var}(\bm{z}^{d})}\right)$. The invariance ($s$) component minimizes the distance between the anchor and positive to learn invariance to perturbations created through data-augmentation, such that $s\left(\bm{Z},\bm{Z}^{\prime}\right)=\frac{1}{B}\sum_{i\in\mathcal{B}}\left\|\bm{z}_{i}-\bm{z}_{i}^{\prime}\right\|_{2}^{2}$. The covariance ($c$) component makes the off-diagonal coefficients of the embeddings covariance matrix $C$ to be close to 00 to prevent the dimensions of the embeddings from encoding similar information, such that $c\left(\bm{Z}\right)=\frac{1}{D_{\text{emb}}}\sum_{d=1}^{D_{\text{emb}}}\sum_{\begin{subarray}{c}d^{\prime}=1\\ d^{\prime}\neq d\end{subarray}}^{D_{\text{emb}}}[C(\bm{Z})]_{d,d^{\prime}}^{2}$.
@@ II-E DINO
DINO [13] is a self-distillation framework based on the knowledge distillation paradigm where a student model is trained to predict the output distribution of a teacher model. Following the “multi-crop” strategy [9], a larger set of augmented utterances with different segment lengths is considered, resulting in four small (local) and two large (global) segments. All views are fed through the student, while only global views are fed through the teacher to learn “local-to-global” correspondences. The teacher is updated with an EMA of the student weights following the asymmetrical joint-embedding architecture. To avoid collapse, sharpening and centering are applied to the teacher outputs: centering prevents one dimension from prevailing while sharpening discourages collapse to the uniform distribution.
@@
$\mathcal{L}_{\text{DINO}}=\frac{1}{B}\sum_{i\in\mathcal{B}}\sum_{t=1}^{2}\sum_{\begin{subarray}{c}s=1\\ s\neq t\end{subarray}}^{6}H\bigg{(}\frac{\bm{z}_{i,t}^{\prime}-\bm{c}}{\tau_{\text{t}}},\frac{\bm{z}_{i,s}}{\tau_{\text{s}}}\bigg{)},$ (5)
where $H\left(\bm{a},\bm{b}\right)=-\operatorname{softmax}(\bm{a})\log\left(\operatorname{softmax}(\bm{b})\right)$, $\bm{z}_{i,t}^{\prime}$ is the $t$-th teacher embedding of the $i$-th sample, $\bm{z}_{i,s}$ is the $s$-th student embedding of the $i$-th sample, $\tau_{\text{t}}$ is the teacher temperature, $\tau_{\text{s}}$ is the student temperature, and $\bm{c}$ is a running mean of the teacher embeddings.
Figure 3: SSL training framework for SV with SSPS-Clustering. Compared to the standard framework, the positive is substituted by a pseudo-positive, which is retrieved from a memory queue based on the anchor’s clustering assignment. Clustering is performed at the beginning of each epoch on reference representations derived from longer and non-augmented audio segments.
@@ III SSPS: Self-Supervised Positive Sampling
Different studies have demonstrated theoretically that the primary factors influencing the quality of SSL representations lie in how the positives are defined [33]. This incorporated “a priori” knowledge helps model class distribution, which is fundamental to learning meaningful representations without labels and human supervision.
@@ III-A Motivation
SSL models trained for SV are prone to learning channel-related information, as anchor-positive pairs are constructed from segments of the same audio utterance, thereby from the same recording. This effect is amplified by training datasets collected ”in the wild”, which contain speech segments corrupted with various channel effects and other factors.
The standard strategy to overcome this bias is to rely on data-augmentation techniques. In SR, the challenge is designing transformations that replicate various acoustic conditions to prevent the model from relying on channel characteristics. However, conventional audio augmentation techniques, such as adding background noise and reverberation, fall short of limiting the impact of the same-utterance positive sampling. These methods do not appear to be sufficient for simulating the diversity among samples from a given same speaker class. Moreover, data-augmentation has inherent limitations since it highly depends on the input data and the downstream task.
Therefore, SSL frameworks lack the ability to match two utterances with different acoustic conditions to the same speaker identity. This represents the primary bottleneck to achieving the performance of supervised systems, which inherently address this challenge using human-annotated labels.
@@ III-B Method
Self-Supervised Positive Sampling (SSPS) is proposed to sample more relevant positives, denoted as pseudo-positives, using the acquired knowledge of the SSL model.
SR datasets, such as VoxCeleb, consist of multiple recordings (or sessions) representing several utterances from a single speaker, and each speaker has multiple recordings. Based on this observation and the fact that SSL representations encode channel information, the following assumption is formulated: “SSL same-utterance positive sampling group representations of the same recordings, sharing similar channel information, before modeling speaker identities”.
Suppose that representations $\bm{y}_{1}$ is from speaker $S$ and recording $R$, $\bm{y}_{2}$ is from the same recording but a distinct utterance, $\bm{y}_{3}$ is from the same speaker but a different recording, and $\bm{y}_{4}$ is from a different speaker. In this case, the assumption can be expressed as $\operatorname{sim}\left(\bm{y}_{1},\bm{y}_{2}\right)\geq\operatorname{sim}\left(\bm{y}_{1},\bm{y}_{3}\right)\geq\operatorname{sim}\left(\bm{y}_{1},\bm{y}_{4}\right)$, where $\operatorname{sim}(\bm{u},\bm{v})$ is the cosine similarity between two vectors. This suggests that the latent space is first organized by groups of utterances from the same recording and then by groups of utterances from the same speaker.
SSPS aims to sample accurate pseudo-positives from the same speakers but from different recordings. This approach falls within the family of bootstrapped techniques, as the selection of pseudo-positives relies entirely on the model’s own evolving representations, without using any labels, and improves over iterations (i.e., enhanced representations lead to better sampling and vice versa). This method can be implemented into any SSL framework and introduces a novel axis of freedom, namely, the design of the positive sampling algorithm. It extends SSL beyond tweaking the architecture of the model, the data-augmentation, or the objective function.
Positive sampling and data-augmentation are complementary strategies that serve the common goal of generating meaningful anchor-positive pairs to shape the underlying speaker distribution in the latent space. Data-augmentation is crucial when employing the standard SSL same-utterance positive sampling, as both the anchor and the positive originate from the same recording and would otherwise have very similar acoustic conditions. However, since SSPS creates anchor-positive pairs from different recordings of the same speaker, it inherently introduces greater variability. It thus has the notable property of being less reliant on data-augmentation, particularly when the training data exhibits diverse acoustic conditions. This property is demonstrated through experiments presented at the end of the article.
@@ III-C Framework
Given a training utterance $u_{i}$ with $i\in\mathcal{B}$, from which the anchor is sampled, let $pos(i)$ denote the index of an utterance $u_{pos(i)}$ used to sample the positive. Standard SSL creates the positive from the same utterance by setting $pos(i)=i$ (Figure LABEL:sub@fig:ssps_overview_ssl) and applying random data-augmentation. In contrast, SSPS selects a pseudo-positive from a different utterance in latent space, such that $pos(i)\neq i$ (Figure LABEL:sub@fig:ssps_overview_ssps).
The proposed framework is built around the following main components: (1) reference representations, extracted from longer unaltered audio segments to obtain a robust latent space of training samples; (2) two memory queues, storing all reference representations and latest positive embeddings throughout training; and (3) pseudo-positive sampling, which determines $pos(i)$ based on the latent space of reference representations and retrieves the corresponding positive embedding from the memory queue.
@@ III-C1 Reference representations
Since SSPS aims to capture the underlying data patterns for identifying speakers and recordings, the anchor and positive representations, $\bm{y}_{i}$ and $\bm{y}_{i}^{\prime}$, are not used to determine $pos(i)$. Instead, a reference segment $\hat{\bm{x}_{i}}$, extracted from $u_{i}$ but without data-augmentation and using a longer audio segment, is introduced. As shown in Fig. 3, the corresponding reference representation $\hat{\bm{y}_{i}}$ is computed during training and used by the SSPS sampling algorithm.
@@ III-C2 Memory queues
SSPS relies on two memory queues, $\hat{\bm{Q}}$ of size $(|\hat{\bm{Q}}|,D_{\text{repr}})$ and $\bm{Q}^{\prime}$ of size $(|\bm{Q}^{\prime}|,D_{\text{emb}})$. $\hat{\bm{Q}}$ is used to store the reference representations $\{\hat{\bm{y}}_{i}\}_{i\in\mathcal{I}}$ and $\bm{Q}^{\prime}$ is used to store the positive embeddings $\{\bm{z}^{\prime}_{i}\}_{i\in\mathcal{I}}$. The former helps the sampling algorithm to determine which samples to use as pseudo-positives (i.e., determining $pos(i)$), while the latter is used to extract the corresponding positive embeddings for the SSL objective function. By default, $|\hat{\bm{Q}}|$ and $|\bm{Q}^{\prime}|$ are set to $N$ to cover all training samples, but $|\bm{Q}^{\prime}|$ can be reduced depending on the setup to limit memory usage. Memory queues are updated dynamically at each training iteration using $\hat{\bm{Y}}$ and $\bm{Z}^{\prime}$, respectively, as illustrated in Fig. 3. For clarity, queues are indexed by training index, meaning $\hat{\bm{q}}_{i}$ and $\bm{q}^{\prime}_{i}$ retrieve the element corresponding to the $i$-th training sample, not the $i$-th position in the queue.
@@ III-C3 Pseudo-positive sampling
After a pre-defined number of training epochs with the standard SSL framework, it is assumed that same-speaker representations are grouped in the latent space but with a large scatter due to the strong representation of channel and recording conditions. This assumption is used to sample a pseudo-positive from the neighborhood of the anchor that does not share the same channel and recording conditions. Pseudo-positive embeddings will be sampled from $\bm{Q}^{\prime}$ if and only if $pos(i)\neq\varnothing$, otherwise, default SSL positive sampling is applied. For most SSL frameworks, the positive embedding of the anchor-positive pair of the $i$-th sample will be $\bm{q}^{\prime}_{pos(i)}$ extracted from $\bm{Q}^{\prime}$, instead of $\bm{z}_{i}^{\prime}$ from the mini-batch, in the objective functions $\mathcal{L}_{\text{SimCLR}}$ (1), $\mathcal{L}_{\text{MoCo}}$ (2), $\mathcal{L}_{\text{SwAV}}$ (3), and $\mathcal{L}_{\text{VICReg}}$ (II-D). For the DINO framework, student outputs are considered to be anchors and teacher outputs to be positives, which implies that the teacher embeddings of the $i$-th sample will be $\bm{q}^{\prime}_{pos(i),1}$ and $\bm{q}^{\prime}_{pos(i),2}$, instead of $\bm{z}_{i,1}^{\prime}$ and $\bm{z}_{i,2}^{\prime}$ from the mini-batch, in the objective function $\mathcal{L}_{\text{DINO}}$ (5).
Note that SSPS can be made symmetric by using $\bm{q}_{pos(i)}$, extracted from a new memory queue $\bm{Q}$ similar to $\bm{Q}^{\prime}$, as a pseudo-positive for $\bm{z}_{i}^{\prime}$. However, this technique was not considered as preliminary experiments demonstrated that it results in equivalent downstream performance.